OpenStudy (anonymous):

(a) find dy, and (b) evaluate dy for the given value of x and dx... y+xy-x=0...x=0...dx=0.01

OpenStudy (anonymous):

\[ dy = \frac{dy}{dx}dx = f'(x)dx \]In this case: \[ y = f(x) = y+xy-x \]

OpenStudy (anonymous):

So your first task is to find \(f'(x)\)

OpenStudy (anonymous):

Oops, I messed up a bit.

OpenStudy (anonymous):

We don't have \(y =f(x)\). Since we have \[ y+xy-x=0 \]

OpenStudy (anonymous):

We have to find \(f'(x)\) using implicit differentiation, I believe.

OpenStudy (anonymous):

Do you need help with implicit differentiation?

OpenStudy (anonymous):

Thanks for your help! I'm just confused about using differentials...using the change in y and change in x as variables....I just don't get it...

OpenStudy (anonymous):

Well, when it comes to differentials, remember that \(\Delta y\) is the actual change in \(y\), where ase \(dy\) is a linear approximation of \(\Delta y\). Obviously if the function isn't a line, it's going to diverge.

OpenStudy (anonymous):



OpenStudy (anonymous):